OpenStudy (anonymous):

A man stands on a cliff that 10m above the sea. He throws a stone vertically upwards with velocity 5ms^-1. The stone eventually lands in the sea, if the air resistance is negligible, what is the time taken for the stone to reach the sea after leaving the man hand. Ans is 14.1 s. Help please!!! :)))

OpenStudy (anonymous):

ans = 0,5 s

OpenStudy (anonymous):

The ans is 14.1 s.

OpenStudy (anonymous):

if u have the ans what do you want ??

OpenStudy (anonymous):

I want the solution.

OpenStudy (anonymous):

time to stop 0 = vo - a t, where a= 9.8 m/s^2 height at top is s = vo t - (1/2) a t^2 distance traveled to sea is 10m + s = (1/2) a T^2 add t + T to get total time